OpenStudy (anonymous):

Two identical objects A and B fall from rest from different heights to the ground. If object B takes twice as long as A to reach the ground, what is the ratio of the heights (A:B) from which the two objects fell? Neglect air resistance.

OpenStudy (anonymous):

D = Vit + 1/2at^2 Falling from rest would remove Vit (Initial Velocity * time) D = 1/2at^2 D/t^2 = 1/2a (in this case a= acceleration = gravity = 9.8 m/s^2 D/t^2 = 4.9 m/s^2 D = 4.9 m/s^2 * t^2 For object A, lets imagine t = 1 and for object B, lets imagine t = 2 for a: D = 4.9 m/s^2 * 1s^2 D = 4.9 m/s^2 * 1s^2 D = 4.9m for b: D = 4.9 m/s^2 * 2s^2 D = 4.9 m/s^2 * 4s^2 D = 19.6m 4.9m/19.6m = .25 Ratio a/b: .5

OpenStudy (anonymous):

1.1:4 2. 1:2 3.1:8 4. 1:squareroot 2

OpenStudy (roadjester):

@greenleaf800073 you're missing a 2

OpenStudy (anonymous):

@k1112 It would be A, 1:4

OpenStudy (roadjester):

4.9m/19.6m = .25 Ratio a/b: .\(\color {red}{\bf 2}\)5